Murgatroyd Co. purchased equipment on January 1, 2016, for $520,000, estimating a four-year useful life and no residual value. In 2016 and 2017, Murgatroyd depreciated the asset using the sum-of-years'-digits method. In 2018, Murgatroyd changed to straight-line depreciation for this equipment. What depreciation would Murgatroyd record for the year 2018 on this equipment? (Do not round your depreciation rate.)

Multiple Choice

  • $ 78,000.

  • $130,000.

  • $156,000.

  • None of these answer choices are correct.
Answer #1

Answer :78,000

Sum of the years digit method :

Totla Digits = 4+3+2+1 = 10

Depreciation for 2016 = 520000*4/10 =208,000

Deprciation for 2017 = 520000*3/10= 156,000

Total depreciation =208,000+156,000 = $364,000

Book value (2018) = 520,000-364,000 = 156,000

Depreication Under Straightline method = 156,000 /2 =78,000 (Answer)